OpenStudy (anonymous):

What is the area of a square with sides of lengths 2x+1 and x+2? (A) 1 (B) 9 (C) 16 (D) 25 (E) It cannot be determined from the information given.

Vocaloid (vocaloid):

all sides of a square are equal so 2x + 1 = x +2 solve for x

OpenStudy (anonymous):

3

Vocaloid (vocaloid):

not quite, 2x + 1 = x + 2 subtract x from each side first, what do we get?

OpenStudy (anonymous):

x+1=2

Vocaloid (vocaloid):

right, now subtract 1 from each side

OpenStudy (anonymous):

x=1

OpenStudy (hybrik):

yes

Vocaloid (vocaloid):

good, so x = 1 now, we know that one of the sides is x+2 if x = 1, then (x+2)^2 = ?

OpenStudy (hybrik):

now plug in one into X

OpenStudy (hybrik):

(2(1)+1)(1+2)

OpenStudy (hybrik):

(2+1)(1+2) = (3)(3) = A

OpenStudy (anonymous):

9

OpenStudy (hybrik):

Correct!

OpenStudy (hybrik):

9[B]